An ungranulated thermometer reads 2.0 and 12.9at ice andsteam point respectively.calculate the true temperature in kelvin when the thermometer reads 5.0

1
Expert's answer
2022-02-07T08:42:38-0500

We use the slope-intercept method to find the temperature in Kelvin when T' = 5.0:


"m=\frac{T_2-T_1}{T'_2-T'_1}=\frac{{373.15\,K}-{273.15\,K}}{12.9-2.0}=\frac{100}{10.9}\,K
\\ \text{We have now: } T-T_1=m(T'-T'_1) 
\\ \therefore T=T_1+m(T'-T'_1)
\\ \text{We have to substitute T'}_1
\\ \text{ to find the temperature T in Kelvin:}
\\ T={273.15\,K}+[\frac{100}{10.9}\,K](5.0-2.0)
\\ T={273.15\,K}+[\frac{300}{10.9}\,K]= 300.67\,K"


In conclusion, the temperature in Kelvin when the thermometer reads 5.0 is T = 300.67 K.